OpenStudy (anonymous):

Which statements are true about the value of the digits in the number 9,999,999?

OpenStudy (welshfella):

starting from the right that 9 = 9 * 10^0 = 9*1 = 9 second = 9 * 10^1 etc This is the base10 number system which is used in evreyday life.

OpenStudy (welshfella):

so can you tell me the value of the first digit on the left is?

OpenStudy (anonymous):

i cant understand how your doing it can you put it in a question that a fifth grader will understand

OpenStudy (missmeow):

Are there any answers you can choose from?

OpenStudy (welshfella):